# Diffusive hydrodynamics from long-range correlations

### v2

@@ Abstract
which typically has exponentially decaying correlations. The hydrodynamic expansion accounts for corrections in terms of the smoothness of the generalized temperatures $\beta^{i}(x)\simeq\beta^{i}(x_{0})+(x-x_{0})\partial\beta^{i}+\ldots$ around each point $x_{0}$. The distribution (1) does not directly represent these corrections; instead, they follow from relaxation processes via a Kubo linear response formalism. This yields the well-known Navier-Stokes (NS) equations for the dynamics of the expectation values of the local densities $\mathcal{q}_{i}(x,t)=\langle q_{i}(x,t)\rangle$ Landau and Lifshitz (1987); Boldrighini and Suhov (1997); Spohn (1991a); Nardis et al. (2019a); Bulchandani (2024):
$\partial_{t}\mathcal{q}_{i}+A_{i}^{~{}k}\partial_{x}\mathcal{q}_{k}=\frac{1}{2\ell}\partial_{x}(\mathfrak{D}_{i}^{~{}k}\partial_{x}\mathcal{q}_{k}),$ (2)
where the flux Jacobian $A_{i}^{~{}k}=\delta\mathcal{j}_{i}/\delta\mathcal{q}_{k}$ describes ballistic transport, with the flux being a function of the local densities $\mathcal{j}_{i}(\mathcal{q}_{\cdot}(x,t))\equiv\langle j_{i}\rangle_{x,t}$; here the state $\langle\cdot\rangle_{x,t}$ is (1) but with $\beta^{i}$ constant, evaluated at $x,t$. The direct (DC) conductivity $\mathfrak{L}_{ij}=\mathfrak{D}_{i}^{~{}k}C_{k,j}$, where the susceptibility matrix is $C_{ik}=\langle Q_{i}q_{k}(0)\rangle^{c}_{x,t}$, is obtained via the Kubo formula $\mathfrak{L}_{i,j}=\int_{-\infty}^{\infty}{\rm d}t^{\prime}\int{\rm d}x^{\prime}\langle j^{-}_{i}(x^{\prime},t^{\prime})j_{j}(0,0)\rangle^{c}_{x,t}$, where the asymptotic ballistic contribution is subtracted out $j_{i}^{-}=j_{i}-A_{i}^{~{}k}q_{k}$, as stated by the celebrated Einstein relation Kubo (1966).
@@
A description as above holds in arbitrary dimensions, but we focus on one dimension of space. In this case, there is an important caveat. When at least one interacting ballistic mode is present, such as in anharmonic chains or interacting one-dimensional gases, Eq. (1) is typically broken: the matrix $\mathfrak{L}$ has divergent components, and superdiffusion arises, as demonstrated in the framework of non-linear fluctuating hydrodynamics Spohn (2014); Spohn and Stoltz (2015); Doyon (2022a); Mendl and Spohn (2016a, b); Popkov et al. (2015); Chen et al. (2018). In certain systems, such as integrable models, it is finite Spohn (1991b); Nardis et al. (2019a); Gopalakrishnan et al. (2018); De Nardis et al. (2018); Doyon (2022a); Medenjak et al. (2020) (with notable exceptions in isotropic spin chains De Nardis et al. (2020); Ljubotina et al. (2019); Wei et al. (2022); Scheie et al. (2021); Das et al. (2019); De Nardis et al. (2023)). As we will show, finiteness of $\mathfrak{L}$ in one-dimensional systems follows if the hydrodynamic equation has a property called linear degeneracy Ferapontov (1991); Rozdestvenskii and Sidorenko (1967); Liu (1979); generalised hydrodynamics (GHD), for integrable systems, has this property El et al. (2011); Pavlov et al. (2012); Bulchandani (2020); Doyon (2020). This property implies that hydrodynamic modes do not “self-interact”; only interactions between modes with different velocities are present. Self-interaction is the source of super-diffusion, as non-linear hydrodynamics is unstable under small noise in one dimension Kardar et al. (1986). Thus, the current understanding is that in linearly degenerate systems, Eq. (2) is expected to describe the space-time profiles of all one-point averages of local observables $\langle o(x,t)\rangle$.
@@
We verify our new equation through Monte Carlo simulations and first-principle calculations Hübner et al. in the hard rods gas Flicker (1968); Robledo and Rowlinson (1986); Doyon and Spohn (2017); Bagchi et al. (2023). Additionally, we derive the evolution equations for all correlation functions, extending the BMFT results. We therefore demonstrate that the Euler-scale propagation of initial fluctuations holds to all orders in $1/\ell$, and here we propose the full all-order theory.
BMFT and equation for diffusive hydrodynamics.—  The BMFT can be formulated as a theory for fluctuating conserved densities Doyon et al. (2023b). One sets a measure on their initial configurations $q_{i}^{0}(x)$, which are then evolved deterministically according to the Euler equation: $q_{i}(x,0)=q_{i}^{0}(x)$ and $\partial_{t}q_{i}(x,t)+\partial_{x}\mathcal{j}_{i}(q_{\cdot}(x,t))=0$. In linearly degenerate systems, and in particular in GHD, the solution is unique (in particular, no shocks appear Ferapontov (1991); Rozdestvenskii and Sidorenko (1967); Liu (1979); Hübner and Doyon (2024)) and thus $q_{i}(x,t)=q_{i}[q^{0}](x,t)$ is a fixed functional of $q_{\cdot}^{0}(\cdot)$: initial fluctuations are deterministically transported. Further, any observable at $x,t$ is taken to be a fixed function of conserved densities at $x,t$, simply its expectation value in the corresponding maximal entropy state, $a(x,t)\equiv\mathcal{a}(q_{\cdot}(x,t))$. That is, for one-point averages, $\langle a(x,t)\rangle\sim\langle\langle\mathcal{a}(q_{\cdot}(x,t))\rangle\rangle$ with
$\langle\langle\mathcal{a}(q_{\cdot}(x,t))\rangle\rangle:=\int[{\rm d}q^{0}]\,e^{-\ell\mathcal{F}_{\rm ini}[q^{0}]}\mathcal{a}(q_{\cdot}[q^{0}](x,t)),$ (4)
where $-\mathcal{F}_{\rm ini}[q]$ is an entropy functional describing the initial state. In a local GGE (1), one has $-\mathcal{F}_{\rm ini}[q^{0}]=\int{\rm d}x\,s[q_{\cdot}^{0}(x)|\beta^{\cdot}(x)]$, for a relative entropy density $s$ of the fluctuating state regarding the target local GGE (see Doyon et al. (2023b)), which gives $S_{ik}(x,y,0)=\delta(x-y)C_{ik}(x,0)$, the delta-function part of (3). This delta-function part in $S_{ik}(x,y,0)$ represents the finite correlations of conserved densities $q_{i}(x,t)$’s on microscopic lengths, whose exact shapes are lost in this approximate theory. In BMFT, eq. (4) holds at leading order in $\ell^{-1}$, thus one takes the $\ell\to\infty$ limit which concentrates the integral on its saddle point $\mathcal{q}_{i}(x,t)=\langle q_{i}(x,t)\rangle$. This gives $\langle a(x,t)\rangle=\mathcal{a}(\mathcal{q}_{\cdot}(x,t))$ and the Euler hydrodynamic equation $\partial_{t}\mathcal{q}_{i}(x,t)+\partial_{x}\mathcal{j}(\mathcal{q}_{\cdot}(x,t))=0$. Similarly, one obtains the evolution equation for the two-point functions Doyon (2018); Doyon et al. (2023b)
@@
We now claim that eq. (4) correctly describes the microscopic evolution up to, including, order $1/\ell$: namely, by keeping the $1/\ell$ fluctuations around the saddle point, (4) also gives the GHD diffusive terms beyond Euler evolution. In this case, one-point averages $\langle a(x,t)\rangle$ (and higher-point functions such as (3)) receive corrections of order $1/\ell$. In the following, we shall still denote $\mathcal{q}_{i}(x,t)=\langle q_{i}(x,t)\rangle$, including all corrections, with evolution equation $\partial_{t}\mathcal{q}_{i}(x,t)+\partial_{x}\langle j_{i}(x,t)\rangle=0$. We seek the $1/\ell$ corrections to $\langle j_{i}(x,t)\rangle$ (or more generally $\langle a(x,t)\rangle$) which we will give rise to the hydrodynamic equation up to diffusive order.
First, we shall mention that an important property of many physical systems, which we assume here, is that all densities and currents are Parity-Time (PT) symmetric; this implies positivity of the Onsager matrix and thus entropy production under the NS equation Nardis et al. (2019b); Nardis and Doyon (2023). One can show, using PT symmetry, that in a local GGE, local expectations of PT-symmetric observable and their correlations, e.g. $S_{ik}(x,y,0)$, do not receive any $1/\ell$ corrections Nardis and Doyon (2023): $\mathcal{F}_{\rm ini}$ are unchanged. Thus, we still have $\langle a(x,0)\rangle=\mathcal{a}(\mathcal{q}_{\cdot}(x,0))$. We concentrate on such observables.
Second, we should remark the following. Assuming (4) to be accurate up to $1/\ell$, needs particular care. Indeed, delta-correlations $S_{ij}\sim\ell^{-1}C_{ij}\delta(x-y)$ would make any non-linear function of $q_{i}(x)$’s at a given point $x$ (keeping $t$ implicit) simply ill-defined. Physically, random variables should, in fact, be functions of fluid-cell averages, $\epsilon^{-1}\int_{-\epsilon/2}^{\epsilon/2}{\rm d}y\,q(x+y)$, for macroscopic size $\epsilon$ as small as possible. Such functions are well-defined, but delta-correlations give unphysical dependence on the fluid-cell size $\epsilon$. However, these delta-correlations represent fast, microscopic-scale fluctuations which are already encoded within the Euler-scale functions $\mathcal{a}(q_{\cdot})$ used to represent the random variables. Indeed, these functions are averages of the corresponding microscopic observable within maximal entropy states, and such averages involve microscopic fluctuations. Thus, the random variables $\mathcal{a}(q_{\cdot})$ should only be affected by slow, macroscopic-scale fluctuations: those that are of order $1/\ell$. This justifies our main proposal: random variables representing local observables should be regularised as $[\mathcal{a}(q_{\cdot}(x))]^{\rm reg}$, where we perform fluid-cell averaging by avoiding equal coordinates. This, intuitively, only considers potential large-scale fluctuations, which, by the BMFT, are encoded within long-range correlations. That is, we define, for any product of conserved densities, its regularised expression
$[q_{i_{1}}(x)\cdots q_{i_{n}}(x)]^{\rm reg}$ (6)
$=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{n}}\int_{\Lambda_{\epsilon}}{\rm d}y_{1}\cdots{\rm d}y_{n}\,q_{i_{1}}(x+y_{1})\cdots q_{i_{n}}(x+y_{n}),$
where $\Lambda_{\epsilon}$ is $[x-\epsilon/2,x+\epsilon/2]^{n}$ with the condition $y_{i}\neq y_{j}\forall i\neq j$, and write
$\langle a(x,t)\rangle=\langle\langle\mathcal{[}a(q_{\cdot}(x,t))]^{\rm reg}\rangle\rangle,$ (7)
and similarly for higher-point correlation functions.
The next step is to expand (7) via the following cumulant expansion:
$\langle f(q^{\rm reg})\rangle=f(\langle q\rangle)+(1/2)f^{\prime\prime}(\langle q\rangle)\langle[q^{2}]^{\rm reg}\rangle^{\rm c}+\ldots,$ (8)
which gives, for $\mathcal{a}=\mathcal{a}(\mathcal{q}_{\cdot}(x,t))$,
$\langle\langle[\mathcal{a}(q_{\cdot}(x,t))]^{\rm reg}\rangle\rangle=\mathcal{a}+\frac{1}{2}\frac{\delta\mathcal{a}}{\delta\mathcal{q}_{i}\delta\mathcal{q}_{j}}\,\ell^{-1}E^{\rm sym}_{ij}(x,t)+O(\ell^{-2}),$ (9)
@@
where we have used symmetric point-splitting notation of the regular correlation $E^{\rm sym}_{ij}=({E_{ij}(x^{+},x^{-};t)+E_{ij}(x^{-},x^{+};t)})/{2}$. From this, we finally obtain the diffusive-order hydrodynamic equation,
@@
$\partial_{t}E_{ij}+$ $\partial_{x}\big{(}A_{i}^{~{}k}E_{kj}\big{)}+\partial_{y}\big{(}E_{ik}A_{j}^{~{}k}\big{)}$
$=-\delta(x-y)\langle j_{l}^{-},q_{i},q_{j}\rangle C^{lk}\partial_{x}\mathcal{q}_{k},$ (11)
@@
It is possible to write the state at any time $t$ in a form similar to (1), which fully accounts both for long-range two-point functions (3) at leading order in $\ell^{-1}$ (BMFT), and one-point function at leading $O(1)$ and first subleading $O(\ell^{-1})$ order (the regularised representation (7)). Indeed, we show SM that the following “non-local GGE” implements exactly this:
@@
$\ +\int_{x\neq y}{\rm d}x{\rm d}y\,\frac{\gamma^{ij}(x,y)}{2}\delta{q}_{i}(x)\delta{q}_{j}(y)\Big{]},$
@@
In this letter, we go further: we propose that extending the BMFT to a fluctuation theory up to the diffusive scale, give not only the diffusive-scale linear-response equation and anomalous fluctuations, but also, by the cumulant expansion and the regularisation (6), the full hydrodynamic equation.
Validity of the NS equation at early times and in linear response.—  The relation between our fluctuation theory and the standard NS equation (2) can be made more explicit. Starting from a local GGE (1), the early times solutions of our our new hydrodynamic equations (10), ( Diffusive hydrodynamics from long-range correlations) approximately satisfy the Navier-Stoker equation (2). Let us solve the BMFT equation ( Diffusive hydrodynamics from long-range correlations) for long-range correlations explicitly, at early times, with initial condition $E(x,y;0)=0$. Normal modes $n_{j}=R_{j}^{~{}k}q_{k}$, diagonalise the flux Jacobian $A=Rv^{\rm eff}R^{-1}$, where the diagonal matrix $v^{\rm eff}$ contains the modes velocities. One may choose the matrix $R$ such that Spohn (2014) $RCR^{-1}=\mathbf{1}$. By (3) normal-mode correlations take the form $\langle n_{k}(x,t)n_{l}(y,t)\rangle^{\rm c}=\delta(x-y){\bf 1}+\langle n_{k}(x,t)n_{l}(y,t)\rangle^{\rm c}_{\rm lr}$ where the second term is regular. At early times, we only require the delta-terms in ( Diffusive hydrodynamics from long-range correlations) to agree. These are accounted for by jumps, so we make the following ansatz:
$\langle n_{i}(x,t)n_{j}(y,t)\rangle^{\rm c}_{\rm lr}=\frac{1}{2}\Delta_{ij}(x,y)\operatorname{sgn}(x-y)$ (13)
$\qquad+\frac{1}{2}K_{ij}(x,y)\operatorname{sgn}(x-y-(v_{i}^{\rm eff}t-v_{j}^{\rm eff}t)),\quad\mbox{($t$ small)}$)
where $\Delta$ and $K$ are continuous in $x,y$. Applying $x$ any $y$ derivatives on the sign function in the first term on the right-hand side gives the delta function on the right-hand side of ( Diffusive hydrodynamics from long-range correlations), and thus we require
$\Delta_{ij}(x,x)=-\frac{\langle n_{i},n_{j},j^{-}_{k}\rangle C^{kr}\partial\mathcal{q}_{r}}{v_{i}^{\rm eff}-v_{j}^{\rm eff}}.$ (14)
In fact, (14) holds for all times: long-range correlation functions display jumps at equal positions, proportional to the local spatial variation of the densities, and fully determined by the local state. Our numerical analysis shows that the width of the jump is of microscopic (see SM SM , Fig. S1). Eq. (14) makes sense if $\langle n_{i},n_{j},j^{-}_{k}\rangle=0$ whenever $v_{i}^{\rm eff}=v_{j}^{\rm eff}$: this property of three-point coupling degeneracy holds in linearly degenerate systems, as we show SM using hydrodynamic projections Doyon (2022b); Ampelogiannis and Doyon (2023) and non-linear response Doyon (2018, 2022a).
The sign function on the second term in (13) annihilates the left-hand side of ( Diffusive hydrodynamics from long-range correlations) at early times as the rotation matrix $R$ may be considered constant, and normal modes evolve with velocities $v_{i}^{\rm eff}$. We thus see that a front opens up, producing the jump at equal coordinates; the local GGE is unstable, as discontinuities in long-range correlations immediately appear. The initial condition requires $\Delta=-K$, and therefore $E^{\rm sym}_{ij}=\operatorname{sgn}(v_{i}^{\rm eff}-v_{j}^{\rm eff})\frac{\Delta_{ij}}{2}$. Putting this together with (14), we find the expression of the diffusion matrix $\mathfrak{D}_{i}^{~{}s}$ in terms of quadratic-charge projection first obtained in Doyon (2022a) and known to agree with the Kubo formula in integrable systems
$\lim_{t\to 0^{+}}$ $\frac{1}{2}\frac{\delta\mathcal{j}_{i}(q_{\cdot})}{{\delta\mathcal{q}_{r}\delta\mathcal{q}_{k}}}E^{\rm sym}_{rk}$ (15)
$=-\frac{\langle j_{i}^{-},n_{r},n_{l}\rangle\langle j_{k}^{-},n_{r},n_{l}\rangle C^{ks}}{|v_{r}^{\rm eff}-v_{l}^{\rm eff}|}\partial_{x}\mathcal{q}_{s}=-\mathfrak{D}_{i}^{~{}s}\partial_{x}\mathcal{q}_{s}.$
At nonzero times $\Delta$ and $K$ satisfy different evolution equations, thus $\Delta\neq-K$ and corrections to (15) accumulate.
By a similar argument, at linear order, our new diffusive-scale equation reproduces the linearised NS equation, $\partial_{t}\langle q_{i}(x,t)q_{r}(0,0)\rangle^{\rm c}+A_{i}^{~{}k}\partial_{x}\langle q_{k}(x,t)q_{r}(0,0)\rangle^{\rm c}=\frac{1}{2}\mathfrak{D}_{i}^{~{}k}\partial_{x}^{2}\langle q_{k}(x,t)q_{r}(0,0)\rangle^{\rm c}$, within GGEs. This can be done by performing linear response theory on the long-range GGE (12): $\beta\to\beta+\delta\beta$ and $\gamma\to 0+\delta\gamma$, where we note that in a homogeneous GGE, we have $\gamma=0$. As the rotation matrix $R$ can be taken constant, the ansatz (13) gives an exact solution (see SM for the explicit form of $\Delta(x,y)$ and $K(x,y)$). Hence, eq. (15) holds at all times.
@@
Entropy conservation and reversibility.—  Under PT symmetry, the entropy $-\int{\rm d}x\,{\rm Tr}(\rho(x,t)\log\rho(x,t))$ is strictly non-decreasing in $t$ under the NS evolution (2), as a result of the positivity of the Onsager matrix Doyon (2020); Nardis and Doyon (2023). Under the new diffusive-order hydrodynamic equation eqs. (10), ( Diffusive hydrodynamics from long-range correlations), this is no longer the case beyond early times. However, with long-range correlations, a more natural entropy is $-{\rm Tr}(\rho_{\rm LR}(t)\log\rho_{\rm LR}(t))$ using (12). We find that this is constant in time SM : there is no information loss as we consider all relevant correlations, at the required scale, generated under the evolution. Eqs. (10) and ( Diffusive hydrodynamics from long-range correlations) are, in fact, reversible, as they are PT symmetric, in contrast to the NS equation (2). This has important consequences. Indeed, assume that PT symmetry at the Euler scale is that induced by the symmetry of an underlying microscopic dynamics. Then forward microscopic evolution followed by backward evolution at the same time gives the same macrostate, in a way that is robust to errors, as long as such errors keep not only local averages, but also long-range two-point correlation functions unchanged. This is in contrast with generic non-integrable systems, where hydrodynamic shocks can be formed during time evolution; then information inexorably flows to microscopic scales (higher and higher multipoint correlations), and the mere knowledge of the two-point functions is not enough to be able to fully revert time evolution.

### v3

@@ Abstract
which typically has exponentially decaying correlations. Averaging the continuity equation over (1) one obtains $\partial_{t}\langle q_{i}\rangle_{x,t}+\partial_{x}\langle j_{i}\rangle_{x,t}=0$; here the state $\langle\cdot\rangle_{x,t}$ is (1) but with $\beta^{i}$ constant, evaluated at $x,t$. By inverting the relation between $\{\beta_{i}\}$ and $\{\langle q_{i}\rangle\}$, one can express the currents $\langle j_{i}\rangle_{x,t}=\mathcal{j}_{i}(\langle q_{\cdot}\rangle_{x,t})$ as function of local density. Denoting $\mathcal{q}_{i}=\langle q_{i}(x,t)\rangle$ (which at this order of approximation equals $\langle q_{i}\rangle_{x,t}$) the so called Euler hydrodynamic equations become
$\partial_{t}\mathcal{q}_{i}+A_{i}^{~k}\partial_{x}\mathcal{q}_{k}=0,$ (2)
where the flux jacobian $A_{i}^{~k}=\delta\mathcal{j}_{i}/\delta\mathcal{q}_{k}$ describes ballistic transport. They are valid in any many-body systems if the observation length and time scales are sufficiently large.
@@
Since the Euler equations conserve total entropy, in order to understand phenomena such as thermalization it is necessary to consider corrections beyond them. The hydrodynamic expansion accounts for corrections in terms of the smoothness of the generalized temperatures $\beta^{i}(x)\simeq\beta^{i}(x_{0})+(x-x_{0})\partial\beta^{i}+\ldots$ around each point $x_{0}$, meaning that in general $\mathcal{q}_{i}\neq\langle q_{i}\rangle_{x,t}$. The distribution (1) does not directly represent these corrections; instead, they follow from the assumption that fast relaxation occurs, on microscopic time scales, from such a local equilibrium state. This yields the well-known Navier-Stokes (NS) equations Landau and Lifshitz (1987); Boldrighini and Suhov (1997); Spohn (1991a); Nardis et al. (2019a); Bulchandani (2024):
$\partial_{t}\mathcal{q}_{i}+A_{i}^{~k}\partial_{x}\mathcal{q}_{k}=\frac{1}{2\ell}\partial_{x}(\mathfrak{D}_{i}^{~k}\partial_{x}\mathcal{q}_{k}).$ (3)
The direct (DC) conductivity (Onsager matrix) $\mathfrak{L}_{ij}=\mathfrak{D}_{i}^{~k}C_{k,j}$, where the susceptibility matrix is $C_{ik}=\langle Q_{i}q_{k}(0)\rangle^{c}_{x,t}$, is obtained via the Kubo formula $\mathfrak{L}_{i,j}=\int_{-\infty}^{\infty}{\rm d}t^{\prime}\int{\rm d}x^{\prime}\langle j^{-}_{i}(x^{\prime},t^{\prime})j_{j}(0,0)\rangle^{c}_{x,t}$, where the asymptotic ballistic contribution is subtracted out $j_{i}^{-}=j_{i}-A_{i}^{~k}q_{k}$, as stated by the celebrated Einstein relation Kubo (1966). Kubo diffusion – Eq. (3) along with the Kubo formula to evaluate $\mathfrak{D}_{i}^{~k}$ – is well-justified in linear response theory around an equilibrium state, however, in writing (3) one makes the assumption that it remains valid beyond linear response. We will show that this assumption is incorrect, although it is a good approximation at early times from local-equilibrium states (1). Throughout this paper we will assume that all densities and currents are Parity-Time (PT) symmetric; this implies positivity of the Onsager matrix and thus entropy production under the NS equation (3) Nardis et al. (2019b); Nardis and Doyon (2023).
@@
A description as above holds in arbitrary dimensions, but we focus on one dimension of space. In this case, there is an important caveat. When at least one interacting ballistic mode is present, such as in anharmonic chains or interacting one-dimensional gases, Eq. (1) is typically broken: the matrix $\mathfrak{L}$ has divergent components, and superdiffusion arises, as demonstrated in the framework of non-linear fluctuating hydrodynamics Spohn (2014); Spohn and Stoltz (2015); Doyon (2022a); Mendl and Spohn (2016a, b); Popkov et al. (2015); Chen et al. (2018). In certain systems, such as integrable models, it is finite Spohn (1991b); Nardis et al. (2019a); Gopalakrishnan et al. (2018); De Nardis et al. (2018); Doyon (2022a); Medenjak et al. (2020) (with notable exceptions in isotropic spin chains De Nardis et al. (2020); Ljubotina et al. (2019); Wei et al. (2022); Scheie et al. (2021); Das et al. (2019); De Nardis et al. (2023)). As we will show in the SM (see remark after Eq. (S35)), finiteness of $\mathfrak{L}$ in one-dimensional systems follows if the hydrodynamic equation has a property called linear degeneracy Ferapontov (1991); Rozdestvenskii and Sidorenko (1967); Liu (1979); generalised hydrodynamics (GHD), for integrable systems, has this property El et al. (2011); Pavlov et al. (2012); Bulchandani (2020); Doyon (2020). This property implies that hydrodynamic modes do not “self-interact”; only interactions between modes with different velocities are present. Self-interaction is the source of super-diffusion, as non-linear hydrodynamics is unstable under small noise in one dimension Kardar et al. (1986). Thus, the Kubo-diffusion understanding is that in linearly degenerate systems, Eq. (3) is expected to describe the space-time profiles of all one-point averages of local observables $\langle o(x,t)\rangle$; and we will explain how this must be corrected.
@@
We verify our new equation through Monte Carlo simulations and first-principle calculations Hübner et al. in the hard rods gas Flicker (1968); Robledo and Rowlinson (1986); Doyon and Spohn (2017); Bagchi et al. (2023).
BMFT and equation for diffusive hydrodynamics.—  The BMFT can be formulated as a theory for fluctuating conserved densities Doyon et al. (2023b). One sets a measure on their initial configurations $q_{i}^{0}(x)$, which are then evolved deterministically according to the Euler equation: $q_{i}(x,0)=q_{i}^{0}(x)$ and $\partial_{t}q_{i}(x,t)+\partial_{x}\mathcal{j}_{i}(q_{\cdot}(x,t))=0$. In linearly degenerate systems, and in particular in GHD, the solution is unique (in particular, no shocks appear Ferapontov (1991); Rozdestvenskii and Sidorenko (1967); Liu (1979); Hübner and Doyon (2024)) and thus $q_{i}(x,t)=q_{i}[q^{0}](x,t)$ is a fixed functional of $q_{\cdot}^{0}(\cdot)$: initial fluctuations are deterministically transported. Further, any observable at $x,t$ is taken to be a fixed function of conserved densities at $x,t$, simply its expectation value in the corresponding maximal entropy state, $a(x,t)\equiv\mathcal{a}(q_{\cdot}(x,t))$. That is, for one-point averages, $\langle a(x,t)\rangle\sim\langle\hskip-1.69998pt\langle\mathcal{a}(q_{\cdot}(x,t))\rangle\hskip-1.69998pt\rangle$ with
$\langle\hskip-1.69998pt\langle\mathcal{a}(q_{\cdot}(x,t))\rangle\hskip-1.69998pt\rangle:=\int[{\rm d}q^{0}]\,e^{-\ell\mathcal{F}_{\rm ini}[q^{0}]}\mathcal{a}(q_{\cdot}[q^{0}](x,t)),$ (5)
where $\mathcal{F}_{\rm ini}[q]$ is an entropy functional describing the initial state. In a local GGE (1), one has $\mathcal{F}_{\rm ini}[q^{0}]=\int{\rm d}x\,s[q_{\cdot}^{0}(x)|\beta^{\cdot}(x)]$, for a relative entropy density $s$ of the fluctuating state regarding the target local GGE (see Doyon et al. (2023b)), which gives $S_{ik}(x,y,0)=\delta(x-y)C_{ik}(x,0)$, the delta-function part of (4). This delta-function part in $S_{ik}(x,y,0)$ represents the finite correlations of conserved densities $q_{i}(x,t)$’s on microscopic lengths, whose exact shapes are lost in this approximate theory and that are re-scaled to infinity as $\ell\to\infty$ because of the factor $\ell$ in the definition of $S_{ik}(x,y,t)$. In BMFT, Eq. (5) holds at leading order in $\ell^{-1}$, thus one takes the $\ell\to\infty$ limit which concentrates the integral on its saddle point $\mathcal{q}_{i}(x,t)=\langle q_{i}(x,t)\rangle$. This gives $\langle a(x,t)\rangle=\mathcal{a}(\mathcal{q}_{\cdot}(x,t))$ and the Euler hydrodynamic equation $\partial_{t}\mathcal{q}_{i}(x,t)+\partial_{x}\mathcal{j}(\mathcal{q}_{\cdot}(x,t))=0$. Similarly, one obtains the evolution equation for the two-point functions Doyon (2018); Doyon et al. (2023b)
@@
We now claim that Eq. (5) correctly describes the microscopic evolution up to, including, order $1/\ell$: namely, by keeping the $1/\ell$ fluctuations around the saddle point, (5) also gives the GHD diffusive terms beyond Euler evolution. In this case, one-point averages $\langle a(x,t)\rangle$ (and higher-point functions such as (4)) receive corrections of order $1/\ell$. In the following, we shall still denote $\mathcal{q}_{i}(x,t)=\langle q_{i}(x,t)\rangle$, including all corrections, with evolution equation $\partial_{t}\mathcal{q}_{i}(x,t)+\partial_{x}\langle j_{i}(x,t)\rangle=0$. We seek the $1/\ell$ corrections to $\langle j_{i}(x,t)\rangle$ (or more generally $\langle a(x,t)\rangle$) which will give rise to the hydrodynamic equation up to diffusive order.
Since BMFT is formally not a theory on microscopic observables, but a theory on fluid-cell averaged quantities, we need to regularize observables by doing an appropriate fluid cell averaging, denoted by $a]^{\rm{reg}}$, see Eq. (12) in the End Matter. For simplicity at $t=0$ we assume an uncorrelated local GGE state (1): in this state we have $\langle\hskip-1.69998pt\langle\mathcal{[}a(q_{\cdot}(x,0))]^{\rm reg}\rangle\hskip-1.69998pt\rangle=\mathcal{a}(\mathcal{q}_{\cdot}(x,0))$ and if the observable is PT-symmetric $\mathcal{a}(\mathcal{q}_{\cdot}(x,0))=\langle a(x,0)\rangle+O(\ell^{-2})$. We concentrate on such observables.
The next step is to expand (13) via a cumulant expansion, which takes the general form (for expectation $\langle\cdots\rangle$ over some random variable $q$)
$\langle f(q)\rangle=f(\langle q\rangle)+(1/2)f^{\prime\prime}(\langle q\rangle)\langle q^{2}\rangle^{\rm c}+\ldots.$ (7)
[…]
$\langle\hskip-1.69998pt\langle[\mathcal{a}(q_{\cdot}(x,t))]^{\rm reg}\rangle\hskip-1.69998pt\rangle=\mathcal{a}+\frac{1}{2}\frac{\delta\mathcal{a}}{\delta\mathcal{q}_{i}\delta\mathcal{q}_{j}}\,\ell^{-1}E^{\rm sym}_{ij}(x,t)+O(\ell^{-2}),$ (8)
@@
where the regularisation $\cdots]^{\rm reg}$ gave rise to the symmetric point-splitting of the correlation $E^{\rm sym}_{ij}=({E_{ij}(x^{+},x^{-};t)+E_{ij}(x^{-},x^{+};t)})/{2}$. From this, we finally obtain the diffusive-order hydrodynamic equation,
@@
$\partial_{t}(\ell^{-1}E_{ij})+$ $\partial_{x}\big(A_{i}^{~k}\ell^{-1}E_{kj}\big)+\partial_{y}\big(\ell^{-1}E_{ik}A_{j}^{~k}\big)$
$=-\ell^{-1}\delta(x-y)\langle j_{l}^{-},q_{i},q_{j}\rangle C^{lk}\partial_{x}\mathcal{q}_{k},$ (10)
@@
Remark 1: (9) and (Diffusive hydrodynamics from long-range correlations) still contain the length scale $\ell$. In fact, they are invariant under rescaling of $\ell$ while keeping microscopic lengths the same: $\ell^{-1}E_{kj}$ are the physical, long-wavelength, long-range correlations measured just beyond the region of strong microscopic-length correlations, and $\ell^{-1}\delta(x-y)=\delta(x_{\rm micro}-y_{\rm micro})$. With this understanding, one could set $\ell=1$. However, (9) and (Diffusive hydrodynamics from long-range correlations) are asymptotic results, valid, in a coarse-grained way, as an asymptotic expansion in large $x_{\rm micro},t_{\rm micro}$ and large initial state wavelengths.
Remark 2: (9) seems to be reminiscent of the BBGKY approach, where the equation for the one-point function depends on the two-point function. The important difference is that in BBGKY one needs to insert the microscopic two-point function, while $E^{\rm sym}_{rk}$ are coarse-grained long-range correlations.
It is possible to write the state at any time $t$ in a form similar to (1), which fully accounts both for long-range two-point functions (4) at leading order in $\ell^{-1}$ (BMFT), and one-point function at leading $O(1)$ and first subleading $O(\ell^{-1})$ order (in the regularised representation (13)). Indeed, we show SM that the following “non-local GGE” implements this:
@@
$\ +\int_{x\neq y}\hskip-8.5359pt{\rm d}x{\rm d}y\,\frac{\gamma^{ij}(x,y)}{2}\delta{q}_{i}(x)\delta{q}_{j}(y)\Big],$
@@
In this letter, we go further: we propose that extending the BMFT to a fluctuation theory up to the diffusive scale, gives not only the diffusive-scale linear-response equation and anomalous fluctuations, but also, by the cumulant expansion and the regularisation (12), the full hydrodynamic equation.
Relation to Navier-Stokes theory.—  The (linearized version of the) NS equation (3) is well-justified in the linear response regime. In the end matter we derive that (9) and (Diffusive hydrodynamics from long-range correlations) reduces to (3) within linear response theory. Furthermore, we also derive that (3) is valid at infinitesimal macroscopic time $t=0^{+}$. This must hold as the initial state (1) is locally close to equilibrium and thus linear response (3) describes early times (this was proven in the gas of hard rods Boldrighini and Suhov (1997)).
These results show that our new theory reduces to (3) close to equilibrium, but supersedes it for the non-equilibrium states which arise throughout the evolution.
@@
Entropy conservation and reversibility.—  Under PT symmetry, the entropy $\int{\rm d}x\,{\rm Tr}(\rho(x,t)\log\rho(x,t))$ is strictly non-decreasing in $t$ under the NS evolution (3), as a result of the positivity of the Onsager matrix Doyon (2020); Nardis and Doyon (2023). Under the new diffusive-order hydrodynamic equations (9), (Diffusive hydrodynamics from long-range correlations), this is no longer the case beyond early times. However, with long-range correlations, a more natural entropy is ${\rm Tr}(\rho_{\rm LR}(t)\log\rho_{\rm LR}(t))$ using (11). We find that this is constant in time SM : there is no information loss as we consider all relevant correlations, at the required scale, generated under the evolution. Eqs. (9) and (Diffusive hydrodynamics from long-range correlations) are, in fact, reversible, as they are PT symmetric, in contrast to the NS equation (3). This has important consequences. Indeed, assume that PT symmetry at the Euler scale is that induced by the symmetry of an underlying microscopic dynamics. Then forward microscopic evolution followed by backward evolution at the same time gives the same macrostate, in a way that is robust to errors, as long as such errors keep not only local averages, but also long-range two-point correlation functions unchanged. This is in contrast with generic non-integrable systems, where hydrodynamic shocks can be formed during time evolution; then information inexorably flows to microscopic scales (higher and higher multipoint correlations), and the mere knowledge of the two-point functions is not enough to be able to fully revert time evolution.